OpenStudy (anonymous):

find the vertex of this parabola y=5x*2+10x-5

OpenStudy (anonymous):

To find the x-coordinate of the vertex of y = ax^2 + bx + c use x = -b/2a The substitute that value for x in the function rule to find y.

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

x = -10/2*5 = -1 would be the x-coordinate.

OpenStudy (anonymous):

You try the y-coordinate.

OpenStudy (anonymous):

i don't know how to do it

OpenStudy (anonymous):

If you have an x-coordinate of a point, you can just sub the x-value in and solve for y: y = 5(-1)^2 +10(-1) -5 = 5 - 10 -5 = -10 So your vertex is (-1, -10)

OpenStudy (anonymous):

so the answer is (-1,-10)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Here is the graph; notice the vertex at (-1, -10) http://www.wolframalpha.com/input/?i=y%3D5x%5E2%2B10x-5

OpenStudy (anonymous):

thx can you help me with a few more

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

thx should i ask them in these comments or open a new question?

OpenStudy (anonymous):

ask here if you want me to help otherwise other will reply

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

find the axis of semmetry for this parabola: y=5x*2+10x-5 write your answer as an equasion

OpenStudy (anonymous):

We already have the answer from the previous work. The axis of symmetry is a vertical line that runs through the vertex. The equation of all vertical line is x = a, where is a real number. Since we have the x-coordinate of the vertex (which is a in the equation above) the answer is x = -1

OpenStudy (anonymous):

yes!!! thx

OpenStudy (anonymous):

In other words, the axis of symmetry is the x-cordinate of the vertex set equal to x (to make an equation).

OpenStudy (anonymous):

determine the direction that this parabola opens: y=-x*2+6x-8 up or down

OpenStudy (anonymous):

The sign (plus or minus) in front of the x^2 term determines the orientation of the parabola. If the sign is + then the parabola opens upward, if the sign is negative then parabola opens downward. In this case, the sign is neg. so it opens downward.

OpenStudy (anonymous):

thx!!

OpenStudy (anonymous):

welcome. is that it?

OpenStudy (anonymous):

find the vertex of his parabola y=-x*2+6x-8 and no i have about 5 more

OpenStudy (anonymous):

Do you want to try to do this one the way I did the first one, and I can check you?

OpenStudy (anonymous):

no i mean we can later or on Monday if you would like to teach me but i just have to finich 5 more so i can go to lunch i have to get off in like 5 min

OpenStudy (anonymous):

x = -6/2(-1) = 3 y = -3^2 + 6(3) - 8 = -9 + 18 - 8 = 1 Vertex is (3, 1)

OpenStudy (anonymous):

find the axis of symmetry for his parabola y=-x*2+6x-8 write your answer as an equasion

OpenStudy (anonymous):

x = 3

OpenStudy (anonymous):

find the x-intercepts of the parabola with vertex (-4,2) and y-intercept (0,-30) write your answer in this form (x1,y1) (x2,Y2) if necessary round to the nearest hundredth

OpenStudy (anonymous):

There are two versions of the answer, your choice depending on what your teacher wants: y = -2(x + 4)^2 + 2 or y = -2x^2 -16x - 30

OpenStudy (anonymous):

x-intercepts are (-5, 0) and (-3, 0)

OpenStudy (anonymous):

aye you still on I have to restart the test can you help?@gryphon

OpenStudy (anonymous):

I've got about 5-7 minutes...

OpenStudy (anonymous):

ok nvm but can I message you so we can work on Monday?

OpenStudy (anonymous):

I'm probably not online for a large part of the day on Mondays

OpenStudy (anonymous):

when would u be on

OpenStudy (anonymous):

Try tuesday after 11am sporadically throughout the day. I'm happy to help, but there are many that can answer this level of question.

OpenStudy (anonymous):

ok thx and ill defanitly message you hopfully we can ocomplish something later

OpenStudy (anonymous):

ok, later